OpenStudy (anonymous):

If the relationship between the number x of watches that can be sold and the price p (in dollars) per watch is given by 5x + p = 320; what price should be charged, and how many watches should be manufactured per week in order to maximize the profit?

OpenStudy (anonymous):

We have \(p=320-5x\). We want to maximize the prophet which, that's maximizing the prize times the number of sold watches. So, let's \(f(x)=x p(x)=320x-5x^2\). f is maximum at \(320-10x=0 \implies x=32\) watches. The prize for maximum prophet is then \(p=320-5(32)=$160\).

OpenStudy (anonymous):

profit* :)

OpenStudy (anonymous):

I was just watching supernatural :D

OpenStudy (anonymous):

I like it :)